OpenStudy (anonymous):

(sin t+ cos t)^2 / sin t cos t = 2 + sec t csc t

OpenStudy (zzr0ck3r):

foil (sin^2(x) + 2cos(x)sin(x) + sin^2(x))/(sinx*cosx) = (1+2cosxsinx)/cos(x)*sin(x) = 1/(cosxsinx) + 2 = 2+csc(x)sec(x)

OpenStudy (zzr0ck3r):

understand @tbrunson ?

OpenStudy (anonymous):

after you foiled it i got lost